the center of the is one focus of earth's orbit around the sun the lenght of the major axis is 186000000 miles and the foci are 3200000 miles apart what the eccentricity of earth's orbit
Answer by lwsshak3(11628)

the center of the is one focus of earth's orbit around the sun the lenght of the major axis is 186000000 miles and the foci are 3200000 miles apart what the eccentricity of earth's orbit
**
eccentricity=c/a
2a=length of major axis=186*10^6
a=93*10^6
2c=32*10^5
c=16*10^5=1.6*10^6
eccentricity=c/a=(1.6*10^6)/(93*10^6)=1.6/93≈0.0172